OpenStudy (anonymous):

A truck pulls a trailer from a standstill with a constant force. After 2 minutes, the truck is moving at 48 km/hr. If the same truck applies the same constant force to pull a second trailer which is 3 times the mass of the first trailer, how fast will the truck be moving after 2 minutes? A. 8 km/hr B. 16 km/hr C. 24 km/hr D. 64 km/hr

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@ags2658

OpenStudy (anonymous):

b i think

OpenStudy (anonymous):

basically \[48 \div 8 =16\]

OpenStudy (anonymous):

i mean 48 div 3 = 16

OpenStudy (solomonzelman):

The inital velocity for track 1 is assumed to be 0?

OpenStudy (anonymous):

so is it b

OpenStudy (solomonzelman):

I am asking a question about a problem, not giving the answer... but in any case,

OpenStudy (anonymous):

no thats 48

OpenStudy (solomonzelman):

the force is constant, that means the acceleration is constant (b/c you couldn't be possibly losing mass)

OpenStudy (solomonzelman):

So, after two min it has accelerated from v=0 km/h, to v=48 km/h.

OpenStudy (anonymous):

yes you can its holding 3 times the mass

OpenStudy (anonymous):

so it losses spped because he is going at the same spped he did with 1 trailer so now hes holding 3 times the first trailer tat way you have to divide

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle a=\Delta v (m/s)/\Delta t(s)=48/2(m/s^2)=24(m/s^2) }\)

OpenStudy (anonymous):

how

OpenStudy (solomonzelman):

The force is same, but the mass is three times bigger, so that means the acceleration is 3 times slower, for track 2.

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle a_{\rm ~track~2}=24(m/s^2) /3=8(m/s^2) }\)

OpenStudy (anonymous):

so its c

OpenStudy (anonymous):

You can use this equation to figure out the answer: \[F=ma\] Let's say the force of a random object is 10 N, the acceleration is 2 m/s/s and the mass is 5 kg. \[10 N= (5 kg)(2 m/s/s)\] Let's multiply the mass by three: 5 X 3=15 kg Plugging that into our equation, we get: \[(15 kg)(? m/s/s)=10 N\] Dividing both sides by 15, you get: 0.6666667 m/s/s The answer to your problem must be smaller...

OpenStudy (solomonzelman):

my units are off, km/h excuse me

OpenStudy (anonymous):

ok now im confused

OpenStudy (anonymous):

I simply gave you an example. :D

OpenStudy (solomonzelman):

After two minutes, what is the velocity if it accelerates by 8km/h, and initial velocity is 0?

OpenStudy (solomonzelman):

basically.

OpenStudy (anonymous):

I wanted to show you WHY the speed of the second object will be slower. :)

OpenStudy (anonymous):

so b

OpenStudy (solomonzelman):

Wait, my units might be off somewhere I have to check

OpenStudy (anonymous):

oh so a

OpenStudy (solomonzelman):

yes, I am wrong my units are off.

OpenStudy (anonymous):

ok now i get the answer is a,b,c, or d

OpenStudy (anonymous):

i said that bec im cofused

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle a_{~\rm track~1}=\Delta v (km/hr)/\Delta t(min)=48(km/hr)/2min }\) A mass 3 times larger, will be accelerating 3 times slower. (because force is same: \(m_1a_1=m_2a_2\) ) So your acceleration is: \(\large\color{black}{ \displaystyle a_{~\rm track~2}=\Delta v (km/hr)/\Delta t(min)~~\div 3 \\ =16(km/hr)/2min }\)

OpenStudy (anonymous):

HELP ME PLZ IM SO COFUSED AND THIS IS MY LAST QUESTION

OpenStudy (anonymous):

SO B

OpenStudy (anonymous):

Correct! :D

OpenStudy (solomonzelman):

Now, in two minutes what will you velocity be? Velocity = Acceleration • Time \(\large\color{black}{ \displaystyle a_{~\rm track~2}=16 (km/hr)/2(min)~~\times 2min=16(km/hr) }\)

OpenStudy (anonymous):

that was my first answer the whole time

OpenStudy (solomonzelman):

But we need to show why.....

OpenStudy (anonymous):

i did the math you know

OpenStudy (anonymous):

thx guys i made a 91

OpenStudy (anonymous):

Right, like SolomonZelman said. He showed the formulas, and I followed him while putting in my thoughts. If I had made more details, you could have been even MORE confused.

OpenStudy (anonymous):

Do you know which ones were wrong? I would like to know if I did something incorrectly. :D

OpenStudy (anonymous):

alright guys BYE AND THX FOR THE HELP @ags2658 AND @solomanzelman

OpenStudy (anonymous):

@solomonzelman

OpenStudy (anonymous):

BYE

OpenStudy (anonymous):

You are welcome! Bye! :D